calculate the force of attraction between two balls each of mass 1 kg each when their centres are 10 cm apart . given G=6.67x10^-11 Nm^2Kg^-2?
 

Grade:

3 Answers

Sai Pavan
9 years ago
According to universal law of gravitation , F= G m1m/ r here m1= m2= 1 and            r =10-1 m  , therefore F = 6.67* 10-11* 1 / (10-1)=6.67* 10-9
Kailash
6 years ago
Two spheres of masses 2M & M are initially at rest at a distance R apart due to mumal force of attraction they approach each other when they are at separation R/2 the acceleration of the centre of mass of spheres would be what is the answer of the question
Ayan khan
5 years ago
According to newtons law of gravitation f=g m1m2÷r^2 
Now, here, m1=m2 both have mass 1kg and r=10^-1m 
Therefore, f=6.67×10^-11×1÷[10^-1]^2
F=6.67×10^-9N